OpenStudy (anonymous):

what is the area of a square with side leangth of 9

OpenStudy (uri):

square's area formula=L^2 :(9)^2 All sides are equal. So, 81.

OpenStudy (anonymous):

81 (:

OpenStudy (anonymous):

81

OpenStudy (nailpolisheverywhere):

81

OpenStudy (anonymous):

Multiply the side length to the power of 2(9*9). You'll get 81. Area of square = 81

OpenStudy (anonymous):

A=a^2-->A=9^2-->A=81